The talk time (in hours) on a cell phone in a month is approximated by the probability density function.

f(x) = x-5 / 5h for 5 < x < 10, 1/h for 10 less than or equal to x less than or equal to 15, - x-20 / 5h for 15 less than or equal to x less than or equal 20.

Determine the following:

a. h
b. P(X<30) Round the answers to two decimal places.
c. P(X<16.0) Round the answers to two decimal places.
d. x such that P(X<x)=0.95
1
Expert's answer
2020-07-27T18:40:31-0400

Solution:

a. Let's use the normalization condition

"\int^{+\infty}_{-\infty}P(x)dx=1".

Then

"\int ^{10} _{5}(x-5)/(5h)dx+\int ^{15} _{10}(1/h)dx+\int ^{20} _{15}(-(x-20)/(5h))dx"

"=(x-5)^2\/(10h)|^{10}_{5}+x\/h|^{15}_{10}-(x-20)^2\/(10h)|^{20}_{15}"

"=5\/(2h)+5\/h+5\/(2h)=10\/h" .

"10\/h=1" =>


"h=10".

b. "P(X<30)=\int^{30}_{-\infty}P(x)dx=1".


c. "P(X<16)=\int^{16}_{-\infty}P(x)dx="

"=\int ^{10} _{5}(x-5)/(5h)dx+\int ^{15} _{10}(1/h)dx+\int ^{16} _{15}(-(x-20)/(5h))dx="

"=5\/(20)+5\/10-(x-20)^2\/(10h)|^{16}_{15}=1\/4+1\/2-(16\/100-25\/100)=0.84"


So, "P(X<16)=0.84"


d. If "x<15" then "P(X<x)<P(X<15)=\int^{15}_{-\infty}P(x)dx=\int ^{10} _{5}(x-5)/(5h)dx+\int ^{15} _{10}(1/h)dx=1/4+1/2=0.75"

Therefore "x>15".

If "x>20" then "P(X<x)>P(X<20)=1"

Therefore

"x<20" (1)

So, "P(X<x)=\int^{x}_{-\infty}P(t)dt="

"=\int ^{10} _{5}(t-5)/(5h)dt+\int ^{15} _{10}(1/h)dt+\int ^{x} _{15}(-(t-20)/(5h))dt="

"=1\/4+1\/2-(t-20)^2\/(10h)|^{x}_{15}=3\/4-((x-20)^2\/100-25\/100)=1-(x-20)^2\/100"

From the condition "P(X<x)=0.95" we obtain the equation

"1-(x-20)^2\/100=0.95"

"(x-20)^2\/100=0.05"

"(x-20)^2=5"

"x_1=20-\sqrt5"

"x_2=20+\sqrt5" - does not satisfy the condition (1).

So,

"x=20-\sqrt5"


"x\approx17.76"